OpenStudy (anonymous):

Solve the system, using substitution. Show work please! Thanks! x = 4y and 2x + 3y = 44

OpenStudy (anonymous):

x=4y substitute x in 2nd equation 2(4y)+3y=44 8y+3y=44 11y=44 y=44/11 y=4 now substitute y value in one of the equation x=4(4) x=16